1.The magnitude of the restoring force that acts on a 1.0-kg pendulum when it is displaced 2.0° from equilibrium is ___ N. (Give your answer with the correct number of significant digits and do not include units in your answer)


2.The magnitude of the restoring force that acts on a 1.0-kg pendulum when it is displaced 10.0° from equilibrium is ___ N. (Give your answer with the correct number of significant digits and do not include units in your answer)


3.On earth the period of a 1.50-m pendulum is ___s. (Give your answer with the correct number of significant digits and do not include units in your answer)


4.On the hypothetical planet Xeon, a pendulum with a length of 95.0 cm swings with a frequency of 1.50 Hz. The acceleration due to gravity on Xeon is ___ m/s2. (Give your answer with the correct number of significant digits and do not include units in your answer)



1
Expert's answer
2021-07-19T15:38:39-0400

1)


"F=(1)(9.8)\sin{2}=0.34\ N"

2)


"F=(1)(9.8)\sin{10}=1.7\ N"

3)


"T=2\pi\sqrt{\frac{1.5}{9.8}}=6.04\ s"

4)


"\frac{1}{1.5}=2\pi\sqrt{\frac{0.95}{g}}\\g=84.4\frac{m}{s^2}"